OpenStudy (anonymous):

Find the volume of the solid whose base is the region between curve 2(sinx)^.5 and x axis over the x interval [0,pi] whose cross section are squares that are perpendicular to x axis

OpenStudy (anonymous):

Well, if the cross sections are squares, you know the height of the square is the same as the height of the function. So you have A of the square is the height of the function squared. Or:4sin(x). So you want to integrate: \[4 \int\limits_{0}^{\pi}\sin(x) dx\].

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :P